OpenStudy (anonymous):

What is the difference in simplest form? PLease help! medal given! (n^2+3n+2)/(n^2+6b+8) - (2n)/(n+4)

OpenStudy (anonymous):

(n+1)(n+2)/(n+2)(n+4)-(2n)/(n+4) =(n+1)/(n+4)-(2n)/(n+4) =(n+1-2n)/(n+4) =(1-n)/(n+4)